# Matrix-Free Geometric Multigrid Preconditioning Of Combined Newton-GMRES For Solving Phase-Field Fracture With Local Mesh Refinement

### v1

@@ 1 Introduction
Based on prior work [14], a GMRES (generalized minimal residual method) iterative linear solver is employed. This is preconditioned with a matrix-free geometric multigrid (GMG) method. In the matrix-free context, the system matrix is not fully assembled [19] and reduces the memory consumption. At the same time, this limits the choice of available smoothers for the multigrid preconditioner. Here, a Chebyshev-Jacobi smoother is employed as it only requires matrix-vector products and an estimate of the largest eigenvalue. Moreover, the inverse diagonal of the system matrix is required, which however can be obtained from the local assembly without required the entire matrix. With these ingredients a matrix-free solution can be set up. Recent works of matrix-free solvers include problems in finite-strain hyperelasticity [7], phase-field fracture [15, 14], Stokes [16], generalized Stokes [29], fluid-structure interaction [30], discontinuous Galerkin [20], compressible Navier-Stokes equations [10], incompressible Navier-Stokes and Stokes equations [9] as well as sustainable open-source code developments [24, 6], matrix-free implementations on locally refined meshes [23], implementations on graphics processors [21], and performance-portable methods on CPUs and GPUs applied to solid mechanics [5]. The main contribution of this work is that we combine a GMG preconditioner with locally refined meshes and primal-dual active set for the inequality constraint of phase-field fracture using local smoothing [13]. Our overall numerical solver is applied to one numerical test, namely Sneddon’s example that is nowadays considered as a benchmark [26]. The outline of these conference proceedings is as follows. In Section 2 the phase-field problem is stated. Then, in Section 3, the numerical solutions are explained, namely the nonlinear solution via a combined Newton method and the linear solution with GMRES and matrix-free geometric multigrid preconditioning. Finally, in Section 4 a numerical test substantiates our developements.
@@ 2 Problem Formulation: The Phase-Field Model
where $X:Y$ denote the Frobenius product of two vector fields. If there is no subscript provided, $L^{2}$-product over the whole material domain $\Omega$ is meant. Energy minimization problems in the phase-field fracture context (e.g. [4, 22]) usually consist of a displacement variable $u:\Omega\rightarrow\mathbb{R}^{d}$ and a phase-field variable $\varphi:\Omega\rightarrow[0,1]$. The phase-field variable $\varphi$ can be understood as an indicator function: It is defined such that $\varphi=0$ where the material is fully broken and $\varphi=1$ where it is intact. Only allowing a fully broken or a completely intact domain leads to discontinuities, which is treated by the Ambrosio-Tortorelli approximation [1, 2]. With this, we introduce a transition zone where $10<\varphi<$ 1 of width $2l$. In the following $l$ is called the length-scale parameter. Deriving the Euler-Langrange equations from computing directional derivatives, the solution sets on the continuous level are given by $\mathcal{V}\coloneqq H^{1}_{0}(\Omega),\mathcal{K}^{n}\coloneqq\{\psi\in\mathcal{W}\,|\,\psi-\varphi^{n-1}\leq 0\text{ a.e. in }\Omega\}$, where $\mathcal{W}\coloneqq H^{1}(\Omega)$. The function space $\mathcal{K}^{n}$ is a convex set arising from the crack irreversibility constraint $\partial_{t}\varphi\leq 0$. In the quasi-static setting, this translates to $\varphi^{n}\leq\varphi^{n-1}$ with $\varphi^{n}:=\varphi(t_{n})$ and $\varphi^{n-1}:=\varphi(t_{n-1})$. This results in an incremental grid $t_{0},\dots t_{N}$ with the step-size $k_{n}=t_{n}-t_{n-1}$, where $t_{0}$ is the initial configuration and $t_{N}$ the end-time configuration. The Euler-Lagrange equations are then given by [22, 31, 18]:
@@
[…]
The active set constant $c$ can be chosen arbitrarily. However, in other contributions [12, 25, 27], the authors state that it can have an influence on the performance. Following our prior work [18], we set $c$ as
$c=c^{k}\coloneqq 2\left|\frac{\lambda_{h,i}^{k}}{(\varphi_{h,i}^{k}-\varphi_{h,i}^{\operatorname{old}})}\right|.$ (11)
@@
This section describes the numerical methods we use to solve the introduced variational phase-field fracture problem. As discretization, we employ a finite element method with $H^{1}$-conforming bilinear finite elements for both the displacement and the phase-field, as an outer solver, we employ a combined Newton active set method and as an inner linear solver, we utilize a GMRES algorithm together with a geometric multigrid preconditioner. The newly developed code is implemented in a matrix-free framework using the finite element library deal.II [3]. Related work having deal.II as well as basis was done in [14, 15].
@@ 3.1 A Combined Newton Type Algorithm
In combination with an iteration on the active set, we obtain scheme outlined in Algorithm 1.
@@
3:     Determine the active set $\mathcal{A}^{n,k}$ and inactive set $\mathcal{I}^{n,k}$ with (9), (10)
@@
The combined Newton active Fracture With set algorithm has two stopping criteria which have to be fulfilled: On one hand, the Newton residual has to be small enough while on the other hand the active set has to remain unchanged over two consecutive iterations. With the active set convergence, we ensure that we indeed applied the constraints in the right way.
@@ 3.2 A Geometric Multigrid Block Preconditioner For GMRES
In a matrix-free framework, only iterative methods which solely rely on matrix-vector products are applicable as smoothers. We choose a Chebyshev-accelerated polynomial Jacobi smoother. This method requires to precompute the inverse diagonal entries of the system matrix (i.e. the Jacobian) and an estimate for the eigenvalues. The latter can be obtained by employing a conjugate gradient method. On the finer grid, we are mainly interested in smoothing out the highly oscillating error parts, thus it suffices to compute the maximal eigenvalue and then approximate the smoothing range by $[\lambda_{\min},\lambda_{\max}]\approx[\rho\lambda_{\min},1.2\lambda_{\max}]$ with $\rho<1$. On the coarse grid, we then compute both the maximal and the minimal eigenvalue for indeed solving the problem. We apply the preconditioner on the whole block system by performing 1 V-cycle on each of the symmetric positive definite diagonal blocks as follows. On locally refined meshes, we employ local smoothing, where the smoothing on each level is restricted to the interior of the level domains. We want to solve the following preconditioned inner linear system arising from finite element discretization and Newton’s method
$P^{-1}\tilde{G}\delta U=P^{-1}\tilde{R},$ (12)
with
$G\coloneqq\begin{bmatrix}\tilde{G}^{uu}&0\\ \tilde{G}^{u\varphi}&\tilde{G}^{\varphi\varphi}\end{bmatrix}.$ (13)
The preconditioner $P^{-1}$ is then constructed as
$P^{-1}\coloneqq\begin{bmatrix}MG(\tilde{G}^{uu})&0\\ 0&MG(\tilde{G}^{\varphi\varphi})\end{bmatrix},$ (14)
with the mentioned Jacobi-Block smoother applied to each diagonal block. One challenge of the multigrid preconditioner is to transfer the information of possible constraints onto the coarser grids. In our case, we have to deal with 3 different types of constraints: Boundary conditions, active set constraints and hanging node constraints. The latter arise may arise in the case of the adaptive mesh refinement. The constraints can be transferred to coarser grid in a canonical way. For instance, for the active set, those degrees of freedom which are constrained on finer grids are also constrained on coarser grids.
@@ 4 Numerical Test: Sneddon’s Benchmark
$\kappa$ Regularization parameter $10^{-12}$
Number of global refinements $3-6$
@@
The analytical solutions [28] are given by
$\operatorname{TCV}_{\operatorname{ref}}=\frac{2\pi\rho l_{0}^{2}}{E^{\prime}}$
and
$\operatorname{COD}_{\operatorname{ref}}=2\frac{pl}{E^{\prime}}\left(1-\frac{x^{2}}{l_{0}^{2}}\right)^{\frac{1}{2}}.$
Table 2: The TCV-error on five different refinement levels and the average number of GMRES iterations over all incremental steps and Newton steps.
[…]
$0.088$ $30669$ $0.00716698$ $0.0063186$ $18.8\%$ $2.96$
$0.044$ $62037$ $0.00650211$ $0.0063186$ $7.78\%$ $4.8$
$0.022$ $182325$ $0.00621216$ $0.0063186$ $2.99\%$ $16.13$
$0.011$ $653253$ $0.00607532$ $0.0063186$ $0.72\%$ $28.56$
$0.0055$ $2516661$ $0.00600751$ $0.0063186$ $0.4\%$ $50.97$
[…]
@@
In Table 2 and Figure 2, the results of the TCV and the COD on five different refinement levels are depicted and compared to the reference solution. We can clearly observe that the numerical solution tends to the reference solution under grid refinement.

### v2

@@ 1 Introduction
Based on prior work [13], a GMRES (generalized minimal residual method) iterative linear solver is employed. This is preconditioned with a matrix-free geometric multigrid (GMG) method. In the matrix-free context, the system matrix is not fully assembled [18], which reduces the memory consumption. At the same time, this limits the choice of available smoothers for the multigrid preconditioner. Here, a Chebyshev-Jacobi smoother is employed as it only requires matrix-vector products and an estimate of the largest eigenvalue. Moreover, the inverse diagonal of the system matrix is required, which however can be obtained from the local assembly without requiring the assembly of the entire matrix. With these ingredients a matrix-free solution can be set up. Recent works of matrix-free solvers include problems in finite-strain hyperelasticity [5], phase-field fracture [14, 13], Stokes [15], generalized Stokes [28], fluid-structure interaction [29], discontinuous Galerkin [19], compressible Navier-Stokes equations [8], incompressible Navier-Stokes and Stokes equations [7] as well as sustainable open-source code developments [23, 4], matrix-free implementations on locally refined meshes [22], implementations on graphics processors [20], and performance-portable methods on CPUs and GPUs applied to solid mechanics [3]. In the prior work [13], the solver only allowed globally refined meshes. The main contribution of this work is that we combine a GMG preconditioner with locally refined meshes and primal-dual active set for the inequality constraint of phase-field fracture using local smoothing [12]. Our overall numerical solver is applied to one numerical test, namely Sneddon’s example that is nowadays considered as a benchmark [25]. The outline of these conference proceedings is as follows. In Section 2, the phase-field problem is stated. Then, in Section 3, the numerical solutions are explained, namely the nonlinear solution via a combined Newton method and the linear solution with GMRES and matrix-free geometric multigrid preconditioning. Finally, in Section 4 a numerical test substantiates our developements.
@@ 2 Problem Formulation: The Phase-Field Model
where $X:Y$ denote the Frobenius product of two vector fields. If there is no subscript provided, the $L^{2}$-product over the whole material domain $\Omega$ is meant. Energy minimization problems in the phase-field fracture context (e.g. [2, 21]) usually consist of a displacement variable $u:\Omega\rightarrow\mathbb{R}^{d}$ and a phase-field variable $\varphi:\Omega\rightarrow[0,1]$. The phase-field variable $\varphi$ can be understood as an indicator function: It is defined such that $\varphi=0$ where the material is fully broken and $\varphi=1$ where it is intact. Only allowing a fully broken or a completely intact domain leads to discontinuities, which is treated by the Ambrosio-Tortorelli approximation. With this, we introduce a transition zone where $10<\varphi<$ 1 of width $2l$. In the following $l$ is called the length-scale parameter. Deriving the Euler-Langrange equations from computing directional derivatives, the solution sets on the continuous level are given by $\mathcal{V}\coloneqq H^{1}_{0}(\Omega),\mathcal{K}^{n}\coloneqq\{\psi\in\mathcal{W}\,|\,\psi-\varphi^{n-1}\leq 0\text{ a.e. in }\Omega\}$, where $\mathcal{W}\coloneqq H^{1}(\Omega)$. The function space $\mathcal{K}^{n}$ is a convex set arising from the crack irreversibility constraint $\partial_{t}\varphi\leq 0$. In the quasi-static setting, this translates to $\varphi^{n}\leq\varphi^{n-1}$ with $\varphi^{n}:=\varphi(t_{n})$ and $\varphi^{n-1}:=\varphi(t_{n-1})$. This results in an incremental grid $t_{0},\dots t_{N}$ with the step-size $k_{n}=t_{n}-t_{n-1}$, where $t_{0}$ is the initial configuration and $t_{N}$ the end-time configuration. The Euler-Lagrange equations are then given by [21, 30, 17]:
@@
$\lambda^{n}+c(\varphi^{n}-\varphi^{n-1})\leq 0\text{ a.e.\ in }\mathcal{I}^{n},\quad\lambda^{n}+c(\varphi^{n}-\varphi^{n-1})>0\text{ a.e.\ in}\mathcal{A}^{n}.$ (9)
The active set constant $c$ can be chosen arbitrarily. However, in other contributions [11, 24, 26], the authors state that it can have an influence on the performance. Following our prior work [17], we set $c$ as $c=c^{k}\coloneqq 2\left|\lambda_{h,i}^{k}/(\varphi_{h,i}^{k}-\varphi_{h,i}^{\operatorname{old}})\right|.$
@@
In this section, as discretization, we employ a finite element method with $H^{1}$-conforming bilinear finite elements for both the displacement and the phase-field, as an outer solver, we employ a combined Newton active set method and as an inner linear solver, we utilize a GMRES algorithm together with a geometric multigrid preconditioner. The newly developed code is implemented in a matrix-free framework using the finite element library deal.II [1]. Related work having deal.II as well as basis was done in [13, 14].
@@ 3.1 A Combined Newton Type Algorithm
In combination with an iteration on the active set, we obtain the scheme outlined in Algorithm 1.
@@
3:     Determine the active set $\mathcal{A}^{n,k}$ and inactive set $\mathcal{I}^{n,k}$ with (9)
@@
The combined Newton active set algorithm has two stopping criteria which have to be fulfilled: On one hand, the Newton residual has to be small enough while on the other hand the active set has to remain unchanged over two consecutive iterations. With the active set convergence criterion, we ensure that we indeed applied the constraints in the right way. Usually, the first prediction of the active set is very bad. Thus, we provide the final active set from the previous incremental step as initial active set for the next one.
@@ 3.2 A Geometric Multigrid Block Preconditioner For GMRES
In a matrix-free framework, only iterative methods which solely rely on matrix-vector products are applicable as smoothers. We choose a Chebyshev-accelerated polynomial Jacobi smoother. This method requires to precompute the inverse diagonal entries of the system matrix (i.e. the Jacobian) and an estimate for the eigenvalues. The latter can be obtained by employing a conjugate gradient method. On the levels where we are mainly interested in smoothing out the highly oscillating error parts (i.e. all levels $>0$), it suffices to compute the maximal eigenvalue and then approximate the smoothing range by $[\lambda_{\min},\lambda_{\max}]\approx[0.08\lambda_{\max},1.2\lambda_{\max}]$.
On the coarsest grid (level 00), where we solve the problem using the Chebyshev Jacobi smoother, we then compute both the maximal and the minimal eigenvalue for indeed solving the problem. We apply the preconditioner on the whole block system by performing one V-cycle on each of the symmetric positive definite diagonal blocks as follows. To neglect hanging node constraints during preconditioning, we employ a local smoothing approach, where the smoother only acts on the subdomains which are refined to the current level [12]. We want to solve the following preconditioned inner linear system arising from finite element discretization and Newton’s method
$P^{-1}\tilde{G}\delta U=P^{-1}\tilde{R},\;G\coloneqq\begin{bmatrix}\tilde{G}^{uu}&0\\ \tilde{G}^{u\varphi}&\tilde{G}^{\varphi\varphi},\end{bmatrix},\;P^{-1}\coloneqq\begin{bmatrix}MG(\tilde{G}^{uu})&0\\ 0&MG(\tilde{G}^{\varphi\varphi})\end{bmatrix},$ (10)
with the mentioned Jacobi-Block smoother applied to each diagonal block. One challenge of the multigrid preconditioner is to transfer the information of possible constraints onto the coarser grids. In our case, we have to deal with three different types of constraints: boundary conditions, active set constraints and hanging node constraints. The latter are avoided due to local smoothing. The boundary constraints are transported to coarser meshes in a canonical manner. The biggest difficulty is to transfer the active set. Simply categorizing those dofs as active which are active on the next coarser level leads to too much information loss. Thus, we transfer the active set to each level following e.g. [9]: on level $k$, we only categorize a degree of freedom as inactive, if all its direkt neighbours are inactive, otherwise it is categorized as active.
@@ 4 Numerical Test: Sneddon’s Benchmark
$\kappa$ Regularization parameter $10^{-12}h$
Number of global refinements $2$
Number of local refinements $0-8$
@@
The analytical solutions [27] are given by $\operatorname{TCV}_{\operatorname{ref}}=\frac{2\pi\rho l_{0}^{2}}{E^{\prime}}$ and $\operatorname{COD}_{\operatorname{ref}}=2\frac{pl}{E^{\prime}}\left(1-\frac{x^{2}}{l_{0}^{2}}\right)^{\frac{1}{2}}$.
Table 2: Computational results for the Sneddon test with $2$ global and $0-8$ local refinements.
$h$#\##DoFs TCV error ##\##Newton steps $\varnothing$ lin. iter. $\max$#\##lin. iter. Wall-time [s]
$0.7071$ $5043$ $288.00\%$ $19$ $1.95$ $6$ 1.20
$0.3536$ $5745$ $115.00\%$ $21$ $1.71$ $7$ 1.58
$0.1768$ $8445$ $45.90\%$ $20$ $1.80$ $8$ 2.29
$0.0884$ $16953$ $18.70\%$ $24$ $1.70$ $9$ 4.31
$0.0442$ $48321$ $7.67\%$ $25$ $2.08$ $10$ 9.84
$0.0221$ $168609$ $2.88\%$ $29$ $2.69$ $13$ 36.20
$0.0110$ $639537$ $0.62\%$ $29$ $3.21$ $18$ 132.00
$0.0055$ $2502945$ $0.50\%$ $33$ $3.88$ $25$ 633.00
$0.0028$ $9916113$ $1.00\%$ $30$ $5.40$ $40$ 2800.00
@@
In Table 2 and Figure 2, the results of the TCV and the COD on nine different refinement levels are depicted and compared to the reference solution. All the computations were performed on the same machine and on 4 cores. We can clearly observe that the numerical solution tends to the reference solution under grid refinement. The number of linear iterations indicates robustness under mesh refinement. However, there still appear peaks, which assumingly come up due to the bad prediction of the active set in the initial Newton step. Once the active set is well predicted, the number of linear iterations is very stable under (local) mesh refinement.